OpenStudy (binary3i):

if y=|x-2| then dy/dx=?. y and x intersept?

OpenStudy (anonymous):

y'= (x - 2) / |x - 2|

OpenStudy (anonymous):

is the function differentiable?

OpenStudy (anonymous):

let u=x-2 and we know that |u|=sqrt(u^2)

OpenStudy (anonymous):

what about x=2?

OpenStudy (anonymous):

I'm not sure I follow Yifan

OpenStudy (anonymous):

y' = 1; <2,inf] y' = -1 <inf,2], you can write |x-2| = sqrt((x-2)^2) In x = 2 there is no derivative, it is not defined.

OpenStudy (anonymous):

I see now. y' does not exist at x=2. But the derivative is still correct for other values.

OpenStudy (anonymous):

yeah that's right

OpenStudy (anonymous):

Yes, if you try to plot the graph of the derivative you will see a step function in x = 2

OpenStudy (binary3i):

gogind is right. this is a step function ,dy/dx gives the slope and here for x less then 2 the slope is -1 and for greater than 2 it is 1. at x=2 the slope is not difined